### v1

# Ego-to-Exo: Interfacing Third Person Visuals from Egocentric Views in Real-time for Improved ROV Teleoperation

Abstract
Underwater ROVs (Remotely Operated Vehicles) are unmanned submersible vehicles designed for exploring and operating in the depths of the ocean. Despite using high-end cameras, typical teleoperation engines based on first-person (egocentric) views limit a surface operator’s ability to maneuver and navigate the ROV in complex deep-water missions. In this paper, we present an interactive teleoperation interface that (i) offers on-demand “third”-person (exocentric) visuals from past egocentric views, and (ii) facilitates enhanced peripheral information with augmented ROV pose in real-time. We achieve this by integrating a 3D geometry-based Ego-to-Exo view synthesis algorithm into a monocular SLAM system for accurate trajectory estimation. The proposed closed-form solution only uses past egocentric views from the ROV and a SLAM backbone for pose estimation, which makes it portable to existing ROV platforms. Unlike data-driven solutions, it is invariant to applications and waterbody-specific scenes. We validate the geometric accuracy of the proposed framework through extensive experiments of 2-DOF indoor navigation and 6-DOF underwater cave exploration in challenging low-light conditions. We demonstrate the benefits of dynamic Ego-to-Exo view generation and real-time pose rendering for remote ROV teleoperation by following navigation guides such as cavelines inside underwater caves. This new way of interactive ROV teleoperation opens up promising opportunities for future research in underwater telerobotics.
I Introduction
Unmanned submersible vehicles such as ROVs (Remotely Operated Vehicles) play a crucial role in subsea inspection, remote surveillance, and underwater cave exploration [1, 2, 3]. They are particularly useful in inspecting deep-water structures and surveying confined spaces that are beyond the reach of human scuba divers [4, 5]. In a typical mission, ROVs are controlled by human operators from a surface vessel, who are responsible for the safe and efficient maneuvering [6, 7]. The control consoles for teleoperation typically offer real-time data such as the egocentric video feed, pose, velocity, depth, etc. State-of-the-art ROVs can also include autonomous features for atomic tasks such as hovering [8], following navigation guidelines (e.g., cavelines inside underwater caves [9, 10]), gripping objects [11, 12], trajectory estimation for visual servoing, etc.
While the subsea industries and agencies such as NOAA and naval defense teams deploy underwater ROVs with high-end cameras, sonars, and IMUs [2, 13] – safe and efficient teleoperation remains a challenge in adverse visibility conditions and around complex or sensitive structures. The typical first-person feeds from an ROV camera provide very limited information in landmark-deprived underwater scenes. The operators on the surface can only see the egocentric view, often without global or peripheral semantic information around the ROV. Although ROVs can use artificial lights to enhance visibility in low-light scenes, their bright light get reflected and back-scattered by suspended particles directly at the front camera [9], creating glare and large blind spots for the operator. Additionally, the autonomous and semi-autonomous features of ROVs also become erroneous without peripheral positioning in such noisy sensing conditions.
Figure 1: The proposed Ego-to-Exo teleoperation interface is demonstrated for an underwater cave exploration scenario with an ROV. The traditional console interfaces are based on egocentric views (top left), which is limiting and disorienting to a surface operator in noisy low-light conditions. Our method generates on-demand exocentric views from a given EOB (Eye On the Back) viewpoint, i.e., third-person views from behind the ROV. It also projects the ROV point clouds to augment a comprehensive exocentric perspective with real-time ROV pose. As shown, this interactive EOB viewpoint selection option is integrated into a standard BlueROV2 console for a significantly improved teleoperation experience.
In this paper, we address these issues by introducing an AR (augmented reality) inspired ROV teleoperation interface that generates third-person (exocentric) perspectives as well as provides interactive control choices for viewpoint selection. As shown in Fig. 1, the proposed console can generate multiple exocentric views from past egocentric images, with the ROV’s point cloud projected on the images as if it were taken by a third person immediately behind the robot. Our earlier work introduced the idea of EOB (Eye On the Back) visuals [14], which produces a single third-person view from immediately behind the ROV to facilitate better teleoperation. This work develops the AR-based end-to-end pipeline for generating on-demand exocentric perspectives given any EOB viewpoints. Additionally, we integrate the feature for geometrically accurate ROV positioning into those views for an interactive ROV teleoperation experience in challenging underwater applications. With these features, an operator can slide across multiple EOB viewpoints for the best view, see where the ROV is, and get comprehensive peripheral information for safe and efficient teleoperation.
Specifically, we design an efficient Ego-to-Exo (egocentric to exocentric) view generation framework integrated into a monocular visual SLAM system for underwater ROV teleoperation. The proposed Ego-to-Exo algorithm keeps track of the ROV camera poses and exploits a buffer of egocentric views for exocentric view synthesis. We then transform and project a pre-sampled 3D model of the ROV, in the form of a point cloud, into those views to generate realistic augmented visuals with more peripheral information. As seen in Fig. 1, such views offer comprehensive information of the surrounding scene with global semantics. In addition to the views, the integrated SLAM system provides real-time pose and map updates for atomic tasks [15, 16] such as obstacle avoidance, object following, next-best-view planning, etc.
The significance of this work resides in the simplicity of a 3D-geometric formulation of the Ego-to-Exo problem, and its integration into a real-time SLAM system. A monocular vision-only pipeline is chosen to ensure generalized utility and computational efficiency. The proposed Ego-to-Exo solution, EOB viewpoint-based parameterization, and ROV point cloud projection – are carried out by closed-form solutions to ensure real-time performance. As opposed to data-driven approaches for exocentric view synthesis, the proposed framework is invariant to the changes in waterbody style, scene geometry, and application scenario – making it transferable to existing underwater ROV platforms.
We demonstrate the performance and utility of the proposed Ego-to-Exo framework through a series of experiments. First, we prove the geometric validity of our estimation with 2D indoor navigation scenarios utilizing a TurtleBot4. We quantify the geometric validity based on ground plane estimation errors and reprojection errors of known reference points in the scene. Next, we conduct experiments for autonomous underwater cave exploration scenarios where a BlueROV2 is teleoperated inside underwater caves and grotto systems at various geographical locations. Remote teleoperation in such cave systems includes unique challenges such as low visibility, turbid water conditions, moving shadow effects, and hazy blind spots due to artificial lights. In these scenarios, we demonstrate how the generated Ego-to-Exo views, the augmented robot pose, and the updated map can facilitate improved underwater ROV teleoperation.
In particular, we discuss the results and observations for a variety of challenging cases of caveline detection and following in noisy low-light conditions inside multiple underwater caves and grotto systems. We demonstrate that the generated exocentric views embed significantly more information and global context about the scene for safe and efficient robot control. The operators can adjust viewpoints to visualize augmented ROV poses on the map in real-time – facilitating an interactive teleoperation experience.
(a) We used two platforms: a 2D ground robot (TurtleBot4) and an underwater ROV (BlueROV2); their 3D mesh models and sampled point clouds are shown.
(b) The proposed algorithm maintains a buffer of past egocentric views and corresponding ROV poses from the monocular SLAM system. Given an EOB viewpoint $r$, an exocentric view is generated by synthesizing the past egocentric views; the 3D model of the ROV is then rendered onto the view.
Figure 2: The proposed Ego-to-Exo problem formulation is shown; the idea is to generate EOB (Eye On the Back) views with augmented 3D robot’s pose for more informative third-person perspectives to facilitate safer and more efficient ROV teleoperation.
II Related Work & Applications
II-A ROV Deployment & Teleoperation
Safe and efficient robot teleoperation is crucial for marine biology, oceanography, offshore energy, and infrastructure maintenance. Underwater ROVs enable researchers to study deep-sea ecosystems and collect samples from depths beyond the limits of human scuba divers. A few prominent applications are as follows.
Subsea Structure Inspection
Underwater structures operate in challenging remote environments and are susceptible to various hazards with significant economic and environmental consequences. Corrosion/erosion combined with tensile stresses caused by waves and seismic activities lead to cracking, leakage, and malfunction which are regularly inspected by ROVs [1, 17]. ROVs are also deployed for 3D mapping and acoustic profiling of marine structures [18, 19].
Underwater infrastructure maintenance & security
The underwater energy sector [20] and submerged data centers [21, 22] are experiencing rapid growth with increasing demand for regular maintenance and surveillance operations. Once ROVs are teleoperated to the site, they need to perform various autonomous/semi-autonomous tasks to inspect, monitor, and ensure safety and integrity of those infrastructures.
Marine archaeology and cave exploration
Exploring underwater heritage sites [23] and caves [4] are important to unravel archaeological history as well as for water resource management. Intelligent ROV capabilities enable effective mapping, sampling, and exploration of remote sites and challenging overhead environments. Moreover, seasonal monitoring of oceanographic features is performed for photometric studies as well [24, 25].
Beyond these sectors, improved ROV teleoperation interfaces based on our proposed Ego-to-Exo augmented visuals will be useful in coastal habitat restoration and surveying, aquaculture, and search-and-rescue missions as well.
II-B Third-Person Views for ROV Teleoperation
A common issue reported by ROV operators is that using a remote vision platform for teleoperation is like looking through a “soda straw” [26]. This is because the typical ROV controller interfaces are based on egocentric first person camera views – which provide no peripheral vision, resulting in significantly reduced situational awareness [27, 28]. As such, there have been multiple methods proposed to mitigate the issue and provide the operator with more visual data [29].
Two types of commonly used approaches exist for generating an exocentric view for unmanned ground and aerial vehicles. The first method uses an external camera to provide the exocentric view such as using a fixed camera to record the vehicle’s trajectory from a distance [30], mounting the camera on a UAV (unmanned aerial vehicle) that follows above the target [31, 32, 33], attaching the camera to an elevated position on the robot [34], utilizing a follower ROV with a camera to provide the external view [35], or using a fish-eye camera to create a top-down perspective [36]. The second method utilizes integrated sensors on the ROV, such as LiDAR (Light Detection and Ranging) to generate a point cloud of the surrounding environment [37, 38] and use that toward creating an augmented/virtual environment for real-time interfacing and teleoperation [39, 40, 41].
Adapting the aforementioned methods from terrestrial or aerial domains to underwater environments presents inherent challenges. First, sending diver-robot teams [42] is not always an option in complex deep-water missions – which are the major use cases for ROVs. Secondly, installing an external visual system requires significant hardware modifications, e.g., they need to be rugged and pressure-sealed with the rest of the robot, re-calibrated for buoyancy and motion dynamics, and additional tether integration for high-speed exocentric data transfer. Even with all the structural modifications, an external camera will provide only one additional third-person perspective. We attempt to address these issues by synthesizing geometrically accurate exocentric views into a monocular SLAM pipeline given a dynamic viewpoint.
III Problem Formulation
We formulate the Ego-to-Exo problem as a 3D geometric algorithm that involves generating an on-demand EOB (Eye On the Back) view, and then projecting the robot on it for an augmented rendering of the scene; see Fig. 2. The proposed method has the following computational components.
III-A Curating Pose and Image Queue from Monocular SLAM
A monocular SLAM algorithm such as ORB-SLAM3 [43] provides a continuous solution for estimating and tracking camera poses from a sequence of monocular images. We use an ORB-SLAM3-based framework to obtain camera poses of each keyframe location to eventually construct the trajectory map of the teleoperated robot. Our adapter SLAM framework initiates the trajectory estimation process by building a pose buffer of length $n$: ${}^{w}\mathbf{T}\triangleq[\,{}^{w}_{i}\mathbf{T},\,{}^{w}_{i-1}\mathbf{T},\,.\,.\,.,{}^{w}_{i-n+1}\mathbf{T}\,]$, where, ${}^{w}_{i}\mathbf{T}=[{}^{w}_{i}\mathbf{R}_{3\times 3}\,|\,{}^{w}_{i}\mathbf{t}_{3\times 1}]$ denotes camera pose at instance $i$ in global ($world$) frame of reference. The corresponding raw egocentric views $\mathbf{I}$ for each instance is also stored in a queue $\mathbf{I}\triangleq[\,\mathbf{I}_{i},\,\mathbf{I}_{i-1},\,\cdots\,\mathbf{I}_{i-n+1}\,]$. These memory buffers are updated instantaneously as the robot pose changes during teleoperation. We use an empirically tuned threshold to trigger an update only when the pose change is significant to avoid unnecessary updates (when the robot is static).
III-B Generating Exocentric Perspective
Given the pose memory ${}^{w}\mathbf{T}$ and egocentric views $\mathbf{I}$, we formulate the Ego-to-Exo problem of estimating an exocentric view from a reference location $r$, looking toward the robots current location $c$, where $r,c\in[i-n+1,\,i]$ and $r<c$. Typically, $c$ is set to $i$ (most recent available frame), and $r$ remains a free variable with $n$ known samples in memory – to mimic the EOB viewpoint generation.
As shown in Fig. 2(a), we use the ROV point cloud set $\mathbf{P}$ of size $3\times m$ as prior. These $m$ points are transformed from current camera pose ${}^{w}_{c}\mathbf{T}$ to reference camera pose ${}^{w}_{r}\mathbf{T}$ using:
$\mathbf{\tilde{P}}=({}^{w}_{r}\mathbf{R}^{-1}\,{}^{w}_{c}\mathbf{R})\cdot\mathbf{P}+({}^{w}_{c}\mathbf{t}-{}^{w}_{r}\mathbf{t}),$ (1)
where $[{}^{w}_{c}\mathbf{R}\,|\,{}^{w}_{c}\mathbf{t}]$ and $[{}^{w}_{r}\mathbf{R}\,|\,{}^{w}_{r}\mathbf{t}]$ represent the pose for current and reference (target) location in world coordinate, respectively. The transformed point cloud $\mathbf{\tilde{P}}$ is then projected onto the target image plane by using camera intrinsics $\mathbf{K}$ as:
$\begin{bmatrix}\mathbf{u}&\mathbf{v}&\mathbf{1_{m\times 1}}\end{bmatrix}^{T}=\lambda_{1}\,\mathbf{K}\cdot\mathbf{\tilde{P}}.$ (2)
Here, $\mathbf{u}$ and $\mathbf{v}$ vectors denote the pixel locations ($u$,$v$) on image $\mathbf{I}_{r}$ for projection; $\lambda_{1}$ is the scale.
III-C 3D ROV Rendering and Map Update
While the SLAM system constructs a map of the surrounding, the proposed algorithm simultaneously renders the 3D ROV model on the same spatial context. As illustrated in Fig. 2(b), the ROV points $\mathbf{P}$ are transformed to the current camera location and projected based on the relative pose information ${}^{w}_{i}\mathbf{T}$ as follows:
$\mathbf{\tilde{P}}_{map}=\lambda_{2}\,{}^{w}_{i}\mathbf{R}\cdot\mathbf{P}+{}^{w}_{i}\mathbf{t}.$ (3)
Here, $\lambda_{2}$ is the scaling factor for the ROV model. Like all monocular SLAM-based systems [44, 45], our proposed mapping and projection method is up to scale. While the scale can be resolved with additional sensor fusion, the augmented visuals of Eq. 3 are sufficient for teleoperation.
IV Implementation & Evaluation
IV-A Implementation Details
The proposed framework is implemented in ROS Noetic on an Ubuntu 20.04 environment. A ROS node for ORB-SLAM3 is integrated as the monocular SLAM backbone. The buffer queue size $n$ is set to $100$ and the frame separation threshold is set to $0.001$ unit. The ROV point clouds are generated by sampling 3D mesh models of BlueROV2 and TurtleBot4; $10,000$ points are sampled for each model. The same ROS packages are used on both robots with domain-specific empirical tuning for the $\lambda$ parameters.
IV-B Proof of Concept: 2D Indoor Navigation
Experimental setup. The proof-of-concept experiments are conducted with TurtleBot4, a 2D ground robot that can be teleoperated with egocentric views from its front-facing monocular camera. It has only two degrees of freedom (DOF) for linear and angular velocity - which simplifies the motion kinematics for tracking its instantaneous position and orientation. We teleoperate it to collect monocular visual data in office, laboratory, and hallway scenarios. The experiments are designed to validate the proposed algorithm by evaluating ground plane estimation and reprojection errors.
Geometric validation: reprojection error analysis. We first evaluate reprojection errors of known reference points from the generated Ego-to-Exo views and estimated ROV pose. As shown Fig. 3, we use standard checkerboard corners as reference points from egocentric views and then evaluate the reprojection errors for those points from exocentric views. This test is iterated over different sets of past egocentric images, each corresponding to a different EOB distance. As shown in Fig. 3(a), a checkerboard is viewed from different EOB distances (further back into the past) indicated by the parameter $f$. More specifically, $f$ is the number of frames between the current egocentric view and the selected EOB view. The corresponding reprojection error is plotted in Fig. 3(b), which shows how the estimation is accurate for lower values of $f$, and gradually degenerates for $f>100$. This is consistent with our visual observation of the projected ROV point cloud, i.e., it is on the ground plane with accurate orientation based on the SLAM trajectory estimates.
(a) The TurtleBot4 trajectory during teleoperation is shown; here, the $f$ numbers indicate the EOB distance from current to reference frame used for the generated Ego-to-Exo views.
(b) Reprojection errors for reference points (checkerboard corners) are evaluated for different EOB distances ($f$). The estimated ground plane is shown with a cube placed on it for visual observation.
Figure 3: We conduct 2D indoor navigation experiments with a TurtleBot4 to validate the geometric accuracy of our Ego-to-Exo estimation; here, results are visualized for ground plane estimation and reprojection errors of known reference points in the scene.
Geometric validation: ground plane estimation. We also visualize the accuracy of the estimated ground plane in these experiments for qualitative assessments. Fig. 3(b) illustrates two cases: one for $f=70$ with a low reprojection error, and another for $f=260$ with a high error. As seen, the estimated ground plane (and the drawn sample cube) validates the geometric accuracy for $f=70$ case. On the other hand, a misaligned ground plane for $f=260$ case demonstrates the underlying error in pose estimation as well as in the reprojection process. Essentially, the geometric accuracy of our proposed Ego-to-Exo algorithm depends on the trajectory estimation performance of the SLAM system. More importantly, a teleoperator can make the right viewpoint selection in real-time for optimal Ego-to-exo rendering.
Figure 4: The reconstructed 3D map of a cave segment in Devil’s Springs, Florida is shown. The colored points represent the tracked ORB features, whereas the ROV trajectory is shown in green circles. The popup frames show samples of: (a) egocentric images; (b) synthesized exocentric images with rendered ROV pose; (c) the underlying camera pose updates; and (d) an exocentric view of the 3D map.
Computational efficiency. To ensure real-time execution in resource-constrained edge devices onboard standard ROV platforms, we analyze the computational complexity of the Ego-to-Exo algorithm for different configurations. Table I shows the memory requirement of our algorithm for different choices of buffer size. The memory footprint is less than $300$ MB for a buffer size of up to $180$ frames, making it highly efficient. Table II demonstrates that the end-to-end framework maintains a consistent output rate of over $25$ FPS (frames per second), making it suitable for integration in existing teleoperation engines. In fact, the Ego-to-Exo algorithm only adds $6.8\%$ overhead on the SLAM backbone for the global map update.
TABLE I: The memory requirement of Ego-to-Exo algorithm is compared for different buffer length.
  Buffer size (# of frames) $50$ $100$ $200$ $300$ $400$
Memory usage (MB) $65$ $142$ $301$ $455$ $609$
TABLE II: The computational complexity of the proposed Ego-to-Exo framework integrated into a SLAM backbone is reported. Here, the metrics used are: (i) ROS node publish rate of the exocentric image; and (ii) the global map update rate.
  Method SLAM Only SLAM + Ego-to-Exo
  Output image rate $26$ FPS $25.1$ FPS
Map update rate $26$ FPS $25.3$ FPS
IV-C Field Deployment: 3D Underwater Cave Exploration
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV possesses full 6-DOF motion. While the $roll$ motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as a guide [46]. The mission objective is to navigate the ROV $75$-$300$ feet deep inside the cave through its complex structures, and then safely return it to the surface. In addition to evaluating the geometric accuracy and robustness, we consider how informative the generated Ego-to-Exo views are compared to traditional consoles for underwater ROV teleoperation.
Real-time map update and teleoperation. In addition to the Ego-to-Exo view generation and ROV pose rendering, our framework simultaneously updates a 3D map with extracted feature points from the SLAM system. Fig. 4 shows an ROV’s trajectory mapped during an underwater cave mission in Devil’s Springs, Florida. As seen, the generated Ego-to-Exo views embed significantly more peripheral information about the scene. The exocentric view of the ROV pose and its relative distance from cave walls or overhead obstacles are useful to surface operators for obstacle avoidance and efficient decision-making. Additionally, the 3D map shows the ROV’s past trajectory and its current pose which are useful to analyze the mission progress, which is not possible in traditional teleoperation consoles. Such a global view of the trajectory map is also useful during emergency evacuation and recovery. Beyond underwater cave exploration, these features will be crucial in ROV-based subsea surveillance and search-and-rescue operations as well.
Accuracy and robustness. Due to the complex scene geometry and absence of a smooth ground plane inside underwater caves, we adopt a homography estimation approach for performance validation. As shown in Fig. 5, April-Tag [47] corners are used as reference points for reprojection. Specifically, we compute the homography transformation between the egocentric and synthesized exocentric views to visualize the reprojection errors. As Fig. 5 shows, we use a sample 2D logo and project it onto the reference April-Tag surface using the homographic transform. The unskewed planar projection validates the accuracy of the Ego-to-Exo pose estimation and point cloud rendering processes.
Figure 5: A snapshot from our cave exploration scenario is shown: (a) Egocentric view with detected reference points; and (b) The synthesized Ego-to-Exo view with projected ROV point cloud. We use a sample logo and perform the homographic projection on the reference surface to demonstrate the accuracy in pose estimation.
Observations: strengths and limitations. Our experiments reveal some key strengths of the proposed teleoperation framework. First, the generated exocentric views closely resemble the actual EOB views during a smooth trajectory, which is usually the case for subsea exploration and surveying tasks. Second, the buffer memory works as a backup during a temporary failure of the SLAM system, typically observed at turning corners or due to abrupt motion. In such cases, our algorithm retains historical poses from its buffer memory; teleoperators can utilize this for situational awareness to safely anchor or pause the mission until communication is restored. On the other hand, its heavy dependency on the SLAM backbone leads to some inherent limitations. Feature-based monocular SLAM systems often fail in feature-deprived noisy underwater scenes, which leads to inaccurate pose tracking and thus inaccurate Ego-to-Exo view synthesis. Tracking 6-DOF ROV motion from monocular vision is particularly challenging with no additional sensor to recover the scale information [48, 49]. We observe some instances where the estimated ROV pose is incorrectly scaled in the rendering. To address this, multi-sensor fusion-based underwater SLAM backbones [50, 51] can be utilized in more critical applications.
V Improved Underwater ROV Teleoperation
Multiple augmented viewpoints. We validate the utility of our proposed Ego-to-Exo teleoperation interface through further experiments on underwater cave exploration data. Our expedition in cave segments at Devil’s Springs, Florida reveals that when ROVs move slowly against strong currents, extending the viewpoint distance of the generated exocentric views can significantly improve teleoperation. This is achieved by tuning the queue parameters $r$, $c$, and $n$ in the proposed TeleOp interface. We consistently find that exocentric views are more informative, especially for about $5$-$10$ seconds preceding the ROV position during navigation. The multiple preceding views offered by our interface are particularly useful for mapping large structures such as newly discovered cave segments or shipwrecks. As Fig. 6 shows, the past ROV trajectories provide more spatial context in the augmented map, enabling operators to control the ROV efficiently around complex underwater structures.
Figure 6: A demonstration of our adjustable EOB viewpoint feature is shown. Teleoperators can slide across the distance ($f$) and find the best exocentric view, which is $f=200$ in this example.
Efficient teleoperation in complex missions. We also conducted trials inside an underwater cave system at Orange Grove, Florida, and a grotto system at Hudson, Florida with a total of five ROV operators. A BlueROV2 with a regular console was sent down for up to $300$-feet deep penetration inside the caves. When only egocentric views were available, the operators reported that maneuvering the robot by following the caveline was quite challenging because little/no ambient light penetrates inside underwater caves. Despite using powerful lights, problems such as moving shadows and scattered waves create significant blind spots. Consequently, tracking and following the caveline or any other navigation markers [46] without any peripheral view is extremely disorienting to the operator. In some cases, we observe that the cavelines get blended with the texture and features of cave walls in noisy conditions; see Fig. 7. In such scenarios, shifting the viewpoint to exocentric views with more global information allows easier identification of cavelines against the surrounding and overhead cave walls.
As Fig. 7 illustrates, our proposed Ego-to-Exo teleoperation interface generates exocentric views mimicking EOB views, which is helpful to operators for more efficient ROV teleoperation. Additionally, the augmented 3D map displays the robot’s pose, allowing much safer maneuvering of the vehicle to its desired orientation. Post-operation feedback from our ROV operators suggests that the exocentric EOB views are more useful for safe ROV maneuvers. They reported much better teleoperation experience as well as a significantly less cognitive load in conducting complex tasks such as object following and complex structure mapping.
Figure 7: Three challenging scenarios are shown for ROV teleoperation inside underwater caves: (i) caveline is not visible, i.e., blended with the background; (ii) caveline is not in the FOV; and (iii) front camera-light interactions with suspended particles are causing hazy egocentric views. In all cases, our Ego-to-Exo augmented visuals are clearer and more informative to a surface operator.
Safer navigation in hazy low-light conditions. Underwater caves present a unique formation of silt and sediment on their floor that results from erosion over extended periods. The silt is susceptible to disturbance from external factors, such as the motion of underwater ROVs or the turbulence generated by their propellers. Although ROV operators pay close attention to avoid contact with floor and cave-walls, it is often unavoidable due to buoyancy imbalance and strong flow of water. Dislodging the sediments results in cloudy or hazy conditions that obscure visibility. Bright lights from the ROV reflect from these suspended particles and make it even more challenging to capture clear imagery of the surroundings. In such cases, third-person EOB views from behind the ROV offer a clearer and more informative perspective for navigation; see Fig. 7. It improves spatial awareness and helps the operator to safely move away from the sediment formations toward open, accessible areas and avoid obstructing other scuba divers in the process.
VI Conclusion & Future Work
This work presents an AR-based framework to synthesize exocentric camera views from egocentric feed in real-time for improved underwater ROV teleoperation. A pose geometry-based closed-form solution is formulated for the proposed Ego-to-Exo problem and then integrated into a monocular SLAM backbone. The end-to-end pipeline only requires a sequence of past egocentric views to generate an exocentric view with the accurate ROV model projected onto it. The geometric proof-of-concept is validated by ground plane estimation and reprojection error analyses in a series of 2D indoor navigation experiments. Subsequent field experiments are conducted for various challenging ROV teleoperation scenarios inside underwater caves at various locations. These experiments demonstrate that the proposed framework: (i) offers more informative peripheral views, (ii) provides better situational awareness, and (iii) facilitates an interactive ROV teleoperation experience. We are currently exploring more comprehensive multi-sensor fusion-based underwater SLAM backbone such as the SVIn2 [50] for more accurate and robust estimation. We further plan to develop and integrate more interactive features that would serve as a simulation platform for ROV teleoperation and navigation research.
Acknowledgements
This work is supported in part by the NSF grants $2330416$, $1943205$, and $2024741$. The authors would like to acknowledge the help from Woodville Karst Plain Project (WKPP), El Centro Investigador del Sistema Acuífero de Quintana Roo A.C. (CINDAQ), Global Underwater Explorers (GUE), Ricardo Constantino, and Project Baseline in providing access to challenging underwater caves.