### v5

# EgoExo++: Integrating On-demand Exocentric Visuals with 2.5D Ground Surface Estimation for Interactive Teleoperation of Subsea ROVs

Abstract
Underwater ROVs (Remotely Operated Vehicles) are indispensable for subsea exploration and task execution, yet typical teleoperation engines based on egocentric (first-person) video feeds restrict human operators’ field-of-view and limit precise maneuvering in complex, unstructured underwater environments. To address this, we propose EgoExo, a geometry-driven solution integrated into a visual SLAM pipeline that synthesizes on-demand exocentric (third-person) views from egocentric camera feeds. Our proposed framework, EgoExo++, extends beyond 2D exocentric view synthesis (EgoExo) to augment a dense 2.5D ground surface estimation on-the-fly. It simultaneously renders the ROV model onto this reconstructed surface, enhancing semantic perception and depth comprehension. The computations involved are closed-form and rely solely on egocentric views and monocular SLAM estimates, which makes it portable across existing teleoperation engines and robust to varying waterbody characteristics. We validate the geometric accuracy of our approach through extensive experiments of 2-DOF indoor navigation and 6-DOF underwater cave exploration in challenging low-light conditions. Quantitative metrics confirm the reliability of the rendered Exo views, while a user study involving operators demonstrates improved situational awareness, navigation safety, and task efficiency during teleoperation. Furthermore, we highlight the role of EgoExo++ augmented visuals in supporting shared autonomy, operator training, and embodied teleoperation. This new interactive approach to ROV teleoperation presents promising opportunities for future research in subsea telerobotics.
I Introduction
Unmanned submersible vehicles such as ROVs (Remotely Operated Vehicles) play a crucial role in subsea inspection, remote surveillance, and underwater cave exploration [1, 2, 3]. They are particularly useful in inspecting deep-water structures and surveying confined spaces that are beyond the reach of human scuba divers [4, 5]. In a typical mission, ROVs are controlled by human operators from a surface vessel, who are responsible for the safe and efficient maneuvering of the vehicle [6, 7]. The control consoles for teleoperation typically offer real-time data such as the egocentric video feed, pose, velocity, depth, etc. State-of-the-art (SOTA) ROVs can also include autonomous features for atomic tasks such as hovering [8], following navigation guidelines inside underwater caves and overhead structures [9, 10, 11], object manipulation [12, 13], trajectory estimation, etc.
While the subsea industries and agencies such as NOAA and naval defense teams deploy underwater ROVs with high-end cameras, sonars, and IMUs [2, 14] – safe and efficient teleoperation remains a challenge in adverse visibility conditions and around complex or sensitive structures. The typical first-person feeds from an ROV camera provide very limited information in landmark-deprived underwater scenes. The operators on the surface can only see the egocentric view, often without global or peripheral semantic information around the ROV [15, 16]. Although ROVs can use artificial lights to enhance visibility in low-light scenes, their bright light get reflected and back-scattered by suspended particles directly at the front camera, creating glare and large blind spots for the operator [10]. Additionally, the autonomous and semi-autonomous features of ROVs become erroneous without peripheral positioning in such noisy sensing conditions.
In this paper, we address these issues by introducing an AR (augmented reality) inspired ROV teleoperation interface that generates third-person (exocentric) perspectives as well as provides interactive control choices for viewpoint selection. As shown in Figure 1, the proposed console can generate multiple exocentric views from past egocentric images, with a virtual ROV model projected on the images as if it were taken by a third person following the robot. Our early work introduced the idea of EOB (Eye On the Back) visuals [18], envisioning a single third-person view from immediately behind the ROV to facilitate better teleoperation. Our recent work materialized this idea in EgoExo [17], by formalizing an AR-based framework that generates on-demand exocentric imagery from any EOB viewpoint. It also integrated the feature for geometrically accurate ROV positioning into those views. This work further advances this direction of research by introducing EgoExo++, a dynamic 2.5D exocentric visualization – analogous to a bird’s-eye view in terrestrial contexts – that offers an interactive and semantically enriched perspective of the environment. Importantly, our approach is closed-form and solely geometry-driven, ensuring accuracy and real-time efficiency without reliance on data-driven methods or training biases. The envisioned interface supports both fore-aft transitions across multiple EOB views and an interactive, rotatable exocentric perspective, enabling a safer and more informed ROV teleoperation.
Specifically, we introduce an efficient framework for generating egocentric to exocentric (and beyond) visual perspectives integrated into a visual SLAM system for underwater ROV teleoperation. The base EgoExo algorithm keeps track of the ROV camera poses, visual features extracted by SLAM, and exploits a buffer of egocentric views for exocentric view synthesis. We then transform and project a pre-sampled 3D model of the ROV, in the form of a point cloud, into those views to generate realistic augmented visuals with more peripheral information. In parallel, the EgoExo++ pipeline utilizes SLAM-generated feature points to identify ground regions and fuses them into a continuous ground surface where pixel colors are transferred from corresponding image regions. We employ a temporal fuse-and-stack strategy to preserve the historical ground evidence, while the 3D ROV model is projected on the same spatial context. The resulting 2.5D perspective enables operators to interact with the scene using dynamic viewpoints in real-time. As illustrated in Figure 1, these views provide operators with a globally informed and semantically rich understanding of the surrounding environment, supporting interactive control from arbitrary viewpoints. In addition to view synthesis, the SLAM backend delivers real-time updates on camera pose and environmental mapping to better assist with atomic tasks [13, 19, 20] such as obstacle avoidance, object following, next-best-view planning, manipulation, etc.
II Background and Related Work
II-A Third-Person Views for ROV Teleoperation
A common issue reported by ROV operators is that using a remote vision platform for teleoperation is like looking through a “soda straw” [21, 18]. This is because the typical ROV controller interfaces are based on egocentric first person camera views – which provide no peripheral vision, resulting in significantly reduced situational awareness [22, 23]. Researchers have explored both fixed [24, 25] and dynamic [26, 27] viewpoint augmentation methods in contemporary human-machine interface study [28, 29].
Two primary approaches are used for generating exocentric views in unmanned ground and aerial vehicles. The first leverages external cameras to capture the vehicle’s motion from a distance; examples include fixed ground cameras [30], UAV-mounted overhead views [31, 32, 33, 34], elevated on-robot mounts [35], camera-equipped follower ROVs [36], and fisheye lenses for top-down perspectives [37, 38]. The second method utilizes additional onboard sensors, such as LiDAR (Light Detection and Ranging), to generate a point cloud of the surrounding environment [24, 25] and use it to create an augmented/virtual reality for interfacing and teleoperation [39, 38, 40, 41].
Adapting the aforementioned methods from terrestrial or aerial domains to underwater environments presents inherent challenges. Firstly, sending diver-robot teams [42] is not always an option in complex deep-water missions – which are the majority of use cases for ROVs. Secondly, UGVs that utilize past egocentric views [43, 44] primarily rely on GPS-based localization that does not apply to GPS-denied underwater environments. Unlike underwater ROVs, ground vehicles generally operate on a 2D plane with limited pitch and roll variations over rough terrain [45]. Thirdly, installing an external visual system requires significant hardware modifications, e.g., they need to be rugged and pressure-sealed, recalibrated for buoyancy and motion dynamics, and additional tether integration for high-speed data transfer. Even with all the structural modifications, an external camera will provide a single additional third-person perspective.
A range of AR/VR-based teleoperation systems have been developed to enhance operator immersion and augment visual feedback for subsea tasks such as object grasping and manipulation [46, 29, 13, 47], inspection [48, 49], and navigation [50]. These systems commonly support third-person perspectives by embedding the operator within an XR (extended reality) environment that incorporates a digital twin of the ROV [51] and reconstructs the surrounding scene using 3D models. While such immersive interfaces improve situational awareness and control, they often require extensive sensory augmentation (e.g., visual, auditory, haptic) at both the ROV and operator ends [13, 41], which increases hardware demands and complicates real-time deployment.
II-B 2.5D Exocentric View Generation
Generating 2.5D/3D third-person views from front-facing camera is critical for scene understanding, both for human teleoperators and for autonomous vehicles. The challenges lie in extreme viewpoint shift and lack of direct depth cues from monocular inputs. Recent efforts for 2.5D view synthesis can be categorized into two main areas: homography-based geometric projections [52, 53] and generative models using encoder-decoder, adversarial, or transformer-based learning [54, 55].
The geometry-guided CNN proposed by [53] warps frontal images to the top view using a fitted homography matrix. While efficient for structured environments, their approach is limited to the flat-ground assumption and struggles with non-planar surfaces. Zhu et al. [56] introduce an intermediate homography view from generative adversarial network (GAN) to reduce the difficulty of pure geometric transformation. Transformer models such as BEVFormer [54] and BEVDepth [57] integrate temporal or multi-view cues to improve realism in synthesized views at a cost of high computation. Other learning-based approaches fuse multiple camera views or additional sensors (e.g., LiDAR) to generate semantic aerial views [58, 59], diverging from monocular egocentric setups. Unlike these data-driven approaches, we propose a lightweight geometric solution, integrated into a visual SLAM pipeline that offers real-time, interactive third-person perspectives without relying on multi-modal sensory augmentation or additional hardware.
III EgoExo++: Problem Formulation
We formulate the EgoExo++ problem as a 3D geometric algorithm that involves generating an on-demand EOB view, reconstructing the ground surface, and then projecting the ROV model both on 2D and 2.5D context for augmented rendering of the scene; see Figure 2. The proposed method has the following computational components.
III-A Curating ROV Pose and Image Buffer
A monocular SLAM algorithm such as ORB-SLAM3 [60] provides a continuous solution for estimating and tracking camera poses from a sequence of monocular images. We use an ORB-SLAM3-based framework to obtain camera poses of each keyframe location to eventually construct the trajectory map of the teleoperated robot. In our implementation, the SLAM pipeline initiates the trajectory estimation process by building a pose buffer of length : , where, denotes camera pose at instance in global () frame of reference. The corresponding raw egocentric views for each instance is also stored in a queue . These memory buffers are updated instantaneously as the robot pose changes during teleoperation. We use an empirically tuned threshold to trigger an update only when the pose change is significant to avoid unnecessary updates (when the robot is static).
III-B Generating 2D Exo Image
Given the pose memory and egocentric views , we formulate the EgoExo problem of estimating an exocentric view from a reference location , looking toward the robot’s current location , where and . Typically, is set to (most recent available frame), and remains a free variable with known samples in memory – to mimic the EOB viewpoint generation.
We use the ROV point cloud model of size as prior. These points are transformed from current camera pose to reference camera pose using:
where and represent the ROV pose for current and reference (target) location in world coordinate, respectively. The transformed point cloud is then projected onto the target image plane by using camera intrinsics as:
Here, and vectors denote the pixel locations (,) on image for projection; is the scale.
III-C Generating 2.5D Exo Views
In EgoExo++, we reuse the SLAM-generated visual features from the EgoExo pipeline to estimate the ground surface and synthesize a lightweight terrain-aware 2.5D perspective. This process involves four stages: (i) selecting candidate feature points for the ground surface, (ii) fitting the ground plane, (iii) translating texture from image pixels to the estimated surface, and (iv) fusing multiple frames over time for real-time visualization.
Due to the lack of horizon line in open water settings and the uneven geometry of confined underwater spaces (e.g., caves), we incorporate geometric priors based on the camera orientation to initialize the ground region estimation. In the nominal case with zero pitch and roll, the ground remains within the bottom half of the image, separated by a horizontal line at (where is the image height). As the camera pitches downward, this line shifts upward, since a larger portion of the ground comes within the camera’s FOV, and vice versa. A camera roll rotates this dividing line accordingly. By computing this orientation-adjusted imaginary horizon from the known camera pose, we restrict candidates to points that fall within the “ground side” of the image. This prior ensures that no 3D point projecting above the horizon (e.g., from cave walls and ceiling) is selected as ground.
Let be the set of SLAM feature points in the world frame, associated with camera pose at time instance . After imposing the geometric prior and pre-selecting candidate points, we fit a plane via RANSAC [61]:
where is an inlier loss with threshold . To enforce stability, we apply a prior that constrains the plane normal within an angle of the expected vertical direction ( in camera frame). Given the plane and a reference anchor (closest point from camera center to ), we define an orthonormal basis on the plane. Each 3D point is expressed in local coordinates as:
A rectangular grid is constructed on the ground plane, and the sparse heights are interpolated to obtain a smooth elevation field . Each grid vertex
is then reprojected to the image using intrinsics :
Image colors are sampled (bilinear interpolation) to texture the grid, producing a dense D ground surface. To extend the ground beyond a single camera frame, all historical ground patches are accumulated in the global frame. Patches are merged using voxel decimation and Delaunay triangulation [62] to avoid redundancy while preserving continuity. The fused mesh forms a D exocentric perspective with surface geometry and realistic coloring consistent with the egocentric imagery.
III-D ROV Model Rendering and Scene Update
While the SLAM system constructs a sparse map of the surroundings, the proposed algorithm simultaneously renders the 3D ROV point cloud (or mesh for EgoExo++) on the same spatial context. The ROV points are transformed to the current camera location and projected based on the relative pose information as follows:
Here, is the scaling factor for the ROV model. Note that our mapping and projection method is up to scale, like all monocular SLAM-based systems [63, 64]. While the scale can be resolved with additional sensor fusion, the augmented visuals of Eq. 7 are sufficient for teleoperation.
IV Implementation & Evaluation
IV-A Implementation Details
The framework is implemented using ROS Noetic in an Ubuntu environment, running on an Intel Core i9 processor with GB of RAM. A ROS node for ORB-SLAM3 is integrated as the monocular SLAM backbone. The buffer queue size is set to and the frame separation threshold is set to unit (up to scale). The ROV point clouds are generated by sampling 3D mesh models of BlueROV2 and TurtleBot4; points are sampled for each model. The scaling parameters , , and are empirically tuned once for each test sequence according to the scale of the SLAM-generated map, ensuring that the rendered ROV model appears realistic in size. Note that we adopt the North-East-Down (NED) frame convention used by [65], which is local to the SLAM origin (not aligned with Earth’s North/East).
IV-B Proof of Concept: 2D Indoor Navigation
Experimental setup. The proof-of-concept experiments are conducted with TurtleBot4, a 2D ground robot that can be teleoperated with egocentric views from its front-facing monocular camera. It has only two degrees of freedom (DOF) for linear and angular velocity, which simplifies the motion kinematics for tracking its instantaneous position and orientation. We teleoperate it to collect visual data with a USB camera at p resolution in office, laboratory, and hallway scenarios. The experiments are designed to validate the proposed algorithm by evaluating ground plane estimation and reprojection errors.
Geometric validation: reprojection error analysis. We first evaluate the reprojection errors of known reference points from the generated EgoExo views and the estimated ROV pose. We use standard checkerboard corners as reference points from egocentric views and then evaluate the reprojection errors for those points from exocentric views. This test is iterated over different sets of past egocentric images, each corresponding to a different EOB distance. As shown in Figure 3(a), a checkerboard is viewed from different EOB distances (further back into the past), indicated by the parameter . More specifically, is the number of frames between the current egocentric view and the selected EOB view. The corresponding reprojection error is plotted in Figure 3(b), which shows how the estimation is accurate for lower values of , and gradually degenerates for . This is consistent with our visual observation of the projected ROV point cloud, i.e., it is on the ground plane with accurate orientation based on the SLAM trajectory estimates.
| # Ego | Inlier | Plane | Normal | Altitude |
| Frames | Fraction () | RMSE () | Drift () | Drift () |
Geometric validation: ground plane estimation. We adopt four metrics to evaluate the quality of ground plane estimation: inlier fraction, plane RMSE error, temporal drift in plane normal, and temporal drift in altitude. The inlier fraction for each frame reports the ratio of inliers to total candidate points () after RANSAC fit:
where are the fitted plane parameters, are the candidate 3D points, and is the distance threshold (see Eqn. 3). A higher indicates that the majority of candidate points are consistent with a single ground plane. Subsequently, the point-to-plane distances of inliers are calculated to quantify the fitting residual as root-mean-square error (RMSE):
A lower RMSE reflects a tighter fit around the estimated plane. Next, to assess temporal consistency, we compute the angular difference between consecutive plane normals:
The mean angular drift across frames is reported, where low values indicate temporal stability. Finally, the altitude at instance is computed as the vertical distance of the camera center to the estimated plane:
In the absence of true measurement, the computed (scaled) altitude is not meaningful; however, a low deviation across frames indicates that the synthesized 2.5D ground remains consistent for visualization.
In 2D indoor setup, the robot’s camera is rigidly mounted with negligible roll and pitch variation, so the estimated ground-plane normal is expected to align with the camera’s vertical axis and remain stable across frames. Consequently, the plane inlier fraction should be consistently high, and the residual error should approach zero. The results obtained from several trials in office, laboratory, and hallway scenarios are summarized in Table I. A high inlier fraction and low normal drift across frames confirm the robustness of our approach under such structured conditions; please refer to the next section for further evaluation in unstructured settings.
Figure 3(b) illustrates two representative examples from an office scene: one for with a low reprojection error, and another for with a high error. As seen, the estimated ground plane normal validates the geometric accuracy for the case. On the other hand, a misaligned plane normal for the case demonstrates the underlying error in pose estimation as well as in the reprojection process. Essentially, the geometric accuracy of our proposed algorithm depends on the pose estimation performance of the SLAM system.
Computational efficiency. We analyze the computational complexity of the proposed algorithm for different configurations to ensure real-time execution in resource-constrained edge devices onboard standard ROV platforms. Table II shows the memory requirement of our algorithm for different choices of buffer size. The memory footprint is less than MB for a buffer size of up to frames, making it highly efficient. Table III demonstrates that the EgoExo framework maintains a consistent output rate of over FPS (frames per second). The added computation for ground estimation slightly reduces the scene update rate in EgoExo++, but still maintains over FPS, making it suitable for integration in existing teleoperation engines.
| Buffer (# of Frames) | |||||
| Memory Usage (MB) | |||||
| Method | SLAM | SLAM+EgoExo | SLAM+EgoExo++ |
| Exo Image | FPS | FPS | FPS |
| Scene Update | FPS | FPS | FPS |
IV-C Field Deployment: 3D Underwater Caves
Experimental setup. We extend our experiments to underwater cave exploration scenarios, where the ROV performs full 6-DOF motions. While the motion is limited in the standard BlueROV2s, we consider all 6-DOF for teleoperation with the buoyancy change and pressure imbalance caused by water flow at the cave openings. For remote teleoperation, we consider the scenarios where human operators maneuver an underwater ROV from the surface by following the caveline and other navigation markers as guides [9]. The mission objective is to navigate the ROV - feet deep inside the cave through its complex structures, and then safely return it to the surface. The videos are recorded at p resolution with a GoPro11 camera mounted on BlueROV2 and then compressed to p within the EgoExo framework. In addition to evaluating the geometric accuracy, we consider how informative the generated views are compared to traditional consoles for ROV teleoperation.
Real-time map update and teleoperation. In addition to the exocentric view generation and ROV pose rendering, the EgoExo framework simultaneously updates a sparse map with extracted feature points from the SLAM system. Figure 4 shows an ROV’s trajectory mapped during our trial in an underwater cave in Peacock Springs, Florida. As seen, the generated EgoExo views embed significantly more peripheral information about the scene. The exocentric view of the ROV pose and its relative distance from cave walls or overhead obstacles are useful to surface operators for obstacle avoidance and efficient decision-making. Additionally, the 3D map shows the ROV’s past trajectory and its current pose, which are useful to analyze the mission progress, which is not possible in traditional teleoperation consoles. Such a global view of the trajectory is also useful during emergency evacuation and recovery. Beyond cave exploration, these features will be crucial in ROV-based subsea surveillance and search-and-rescue operations as well.
Validation: homographic projection. Due to the complex scene geometry inside underwater caves, we adopt a homography estimation approach for the performance validation. As shown in Figure 5, April-Tag [66] corners are used as reference points for reprojection. Specifically, we compute the homography transformation between the egocentric and synthesized exocentric views to visualize the reprojection errors. We use a sample 2D logo and project it onto the reference April-Tag surface using the homographic transform. The unskewed planar projection validates the accuracy of the EgoExo pose estimation and point cloud rendering processes.
Validation: ground plane and 2.5D view. The ground plane estimation in the field is assessed following the same method as the indoor validation; Table IV summarizes the results for trials performed at two different cave systems. In addition to the four metrics defined earlier in Eqn. 8- 11, a success rate is also reported, since a valid ground plane cannot always be recovered in unstructured cave scenes. The success rate is defined as the percentage of frames in which a plane can be reliably fitted from the sparse SLAM feature points.
| Field Trials | Peacock Springs | Devil’s Springs |
| Ego Frames | segments | segments |
| Success Rate () | ||
| Inlier Fraction () | ||
| Plane RMSE () | ||
| Normal Drift () | ||
| Altitude Drift () | ||
The results in Table IV show that although trials in Devil’s Springs cave systems have a higher success rate in detecting the ground plane, the plane quality from Peacock Springs cave systems is consistently better. This difference can be attributed to the more complex cave structure and the challenging ROV trajectory executed in the latter case. In Devil’s Springs, several obstacles (e.g., large rocks) appeared directly in front of the ROV, forcing the operator to ascend and maneuver around. The terrain itself had high altitude variations, composed of rocks, boulders, and scattered pebbles, in contrast to the relatively smooth sedimentary floor observed in Peacock Springs. The reconstructed ground map from Peacock Springs shows consistent elevation and orientation (see Figure 4(b)), supporting the quantitative results. More snapshots from the two sites are provided in Figure 6. As seen, the rocky terrain in Devil’s Springs and the resulting jerky vehicle motion led to larger errors in ground plane estimation, greater deviations in the fitted normal, and higher variability in estimated altitude.
Observations: strengths and limitations. Our experiments reveal some key strengths of the proposed teleoperation framework. First, the generated exocentric views closely resemble the actual EOB views during a smooth trajectory, which is usually the case for subsea exploration and surveying tasks. Second, the buffer memory works as a backup during a temporary failure of the SLAM system, typically observed at turning corners or due to abrupt motion. In such cases, our algorithm retains historical poses from its buffer memory; teleoperators can utilize this for situational awareness to safely anchor or pause the mission until communication is restored. On the other hand, its heavy dependency on the SLAM backbone leads to some inherent limitations. Feature-based monocular SLAM systems often fail in feature-deprived, noisy underwater scenes, which leads to inaccurate pose tracking and thus inaccurate EgoExo view synthesis. Tracking 6-DOF ROV motion from monocular vision is particularly challenging with no additional sensor to recover the scale information [67, 68]. We observe some instances where the estimated ROV pose is incorrectly scaled in the rendering. To address this, multi-sensor fusion-based underwater SLAM backbones [69, 70] can be utilized in more critical applications.
| # | Questions | Mean, Std. Dev. |
| 1 | I think that I would like to use this system frequently. | 4.30.6 |
| 2 | I found the system unnecessarily complex. | 2.00.7 |
| 3 | I thought the system was easy to use. | 4.30.4 |
| 4 | I think that I would need the support of a technical person to be able to use this system. | 2.00.6 |
| 5 | I found the various functions in this system were well integrated. | 4.00.8 |
| 6 | I thought there was too much inconsistency in this system. | 2.30.7 |
| 7 | I would imagine that most people would learn to use this system very quickly. | 4.40.5 |
| 8 | I found the system very cumbersome to use. | 2.00.6 |
| 9 | I felt very confident using the system. | 3.70.7 |
| 10 | I needed to learn a lot of things before I could get going with this system. | 1.40.5 |
| 11 | The proposed exocentric view is beneficial for ROV teleoperation. | 4.50.5 |
| 12 | I found the EOB distance tuning feature useful to get the best view. | 4.50.5 |
| 13 | The generated 3D map provides a better understanding of the ROV’s global location and its surroundings. | 4.60.9 |
IV-D Subjective User Study
The user study is conducted with multiple underwater cave exploration data collected during our field trials. A BlueROV2 recorded egocentric video feeds inside the caves at up to -meter penetrations. Later on, human participants, between the ages of - with little/no prior teleoperation experiences, evaluate the ease of operation with our developed console and compare it to traditional consoles. Their feedback is recorded using the System Usability Scale (SUS) [71], with our interface achieving an average SUS score of . We also formulate an independent set of questions on the teleoperator’s preference for the novel features of our method. The individual questions and corresponding scores are presented in Table V. Some key observations from this study are listed below.
1)
The obtained SUS score is fairly above median (score: ) and is considered Good for user experience; it is slightly below the Excellent (score: ) category.
2)
Post-operation feedback from our ROV operators suggests that the exocentric views are more useful for safe ROV maneuvers.
3)
The synthesized 3D map provides a better sense of the ROV’s global location and improves spatial awareness of the teleoperators.
4)
The operators report a significantly lower workload (perceived cognitive load) in conducting complex tasks such as object following and structure mapping.
V Improved Underwater ROV Teleoperation: Strengths, Challenges, and Limitations
Multiple augmented viewpoints. We validate the utility of our proposed EgoExo teleoperation interface through further experiments on underwater cave exploration data. Our expedition in cave segments at Devil’s Springs, Florida, reveals that when ROVs move slowly against strong currents, extending the exocentric viewpoint distance can significantly improve teleoperation. This is achieved by tuning the queue parameters , , and in the proposed TeleOp interface. We consistently find that exocentric views are more informative, especially for about - seconds preceding the ROV position during navigation. The multiple preceding views offered by our interface are particularly useful for mapping large structures such as newly discovered cave segments or shipwrecks [72, 73]. As Figure 6 shows, the synthesized viewpoints provide more spatial context, enabling operators to control the ROV efficiently around complex underwater structures.
2D and 2.5D exocentric view. Our EgoExo pipeline synthesizes third-person views as flat 2D projections of the robot model onto a reference egocentric view from the past. The EgoExo++ advances from this purely image-based rendering to a 2.5D representation by recovering and texturing a dense ground surface in 3D space. This addition provides both altitude awareness and geometric context relative to the terrain, enabling operators to maintain safe clearance above uneven ground [38]. Insights from our prior user study also emphasized the value of adjustable viewpoints, such as bird’s-eye or side perspectives. While earlier EOB viewpoints offered multiple exocentric views, they remained anchored to the robot’s trajectory and fixed reference frames. The EgoExo++ view is no longer tied to a fixed reference image; the virtual viewpoint can be freely adjusted, offering better situational awareness during ROV teleoperation [25].
Efficient teleoperation in complex missions. We conducted extensive field trials across multiple underwater cave systems, including Orange Grove, Devil’s Springs, and Peacock Springs, as well as inside a grotto system in Hudson, Florida. We observe that maneuvering the robot by following the caveline with egocentric views is challenging because little/no ambient light penetrates inside underwater caves. Despite using powerful lights, problems such as moving shadows and scattered waves create significant blind spots [74]. Consequently, tracking and following the caveline or any other navigation markers [9] without any peripheral view is extremely disorienting to the operator. In some cases, we observe that the cavelines get blended with the texture and features of cave walls in noisy conditions; see Figure 7. In such scenarios, shifting the viewpoint to exocentric views allows easier identification of cavelines against the surrounding and overhead cave walls. Additionally, the augmented 3D map displays the robot’s pose, allowing much safer maneuvering of the vehicle to its desired orientation [75].
Safer navigation in hazy low-light conditions. Underwater caves present a unique formation of silt and sediment on their floor that results from erosion over extended periods. The silt is susceptible to disturbance from external factors [76], such as the motion of underwater ROVs or the turbulence generated by their propellers. Although ROV operators pay close attention to avoid contact with the floor and cave walls, it is often unavoidable due to buoyancy imbalance and strong flow of water. Dislodging the sediments results in cloudy or hazy conditions that obscure visibility [77]. Bright lights from the ROV reflect from these suspended particles and make it even more challenging to capture clear imagery of the surroundings. In such cases, third-person EOB views from behind the ROV offer a clearer and more informative perspective for navigation, as shown in Figure 7. It improves spatial awareness and helps the operator to safely move away from the sediment formations toward open, accessible areas and avoid obstructing other scuba divers in the process [78, 9].
Operator-ROV shared autonomy. In subsea teleoperation, collaborative decision-making frameworks split responsibilities so that humans set high-level goals while the vehicle plans and autonomously executes low-level actions [13]. EgoExo++ augmented visuals strengthen this “human-in-the-loop” paradigm by providing an interactive shared scene where operator intent (e.g., safe altitude, keep-out zones) can be directly expressed and the ROV autonomously confirms execution. For instance, the 2.5D view helps the operator determine a safe ground clearance, which the ROV can then autonomously maintain. Additionally, higher-level human-robot interactions can be integrated in the EgoExo++ interface: the operator can draw an intended path directly on the exocentric view [79], then the ROV can plan and follow an optimal path accordingly. Beyond navigation, augmented visuals have high demand in shared telemanipulation tasks such as delicate object grasping [80], valve control [81], artifact collection [46], etc. Our proposed views provide the operator with critical situational cues in such tasks. For instance, a side-view perspective helps position the ROV with respect to the target, then the operator can switch to a close-up egocentric view for precise manipulation [13, 82]. Overall, EgoExo++ serves as a shared perceptual layer: it enhances situational awareness with explicit geometric cues and allows the operator to specify intent more precisely, promising a safer and more effective teleoperation and telemanipulation.
Digital Twins and shadows. A digital shadow creates a virtual replica of the robot and its environment, enabling operators to practice missions, rehearse manipulation tasks, and refine control strategies [83]. While a shadow is a passive replica, a digital twin offers a bidirectional data pipeline and predictive simulation, thereby closing the feedback loop [84]. In this context, EgoExo++ complements DT-based training by providing geometrically consistent exocentric perspectives and interactive Real2Sim 2.5D reconstructions derived from mission SLAM data. During rehearsal, such views allow operators to anticipate spatial challenges, practice navigation in cluttered cave-like terrains, and visualize manipulators reaching the target from third-person perspectives [85]. EgoExo++ views can also be rendered on HMIs, where identical head motions may be mapped to different outcomes depending on the selected visualization mode [86]. For instance, a head tilt in egocentric mode can directly control the ROV body, whereas the same movement in exocentric mode can control the virtual camera viewpoint, with no impact on the ROV. Rehearsing such multi-visual feedback and control mappings in high-fidelity simulator engines will significantly improve operator skills in high-risk, time-critical missions.
Challenges in ground estimation. Our field trials in diverse underwater caves and grotto systems reveal that the irregular and deceptive terrain poses several unique challenges in estimating the ground surface. As illustrated in Figure 8, elevation slope or sharp jumps may fall outside the fitting capability of plane detection algorithms, leading to fragmented or distorted ground estimation. Additionally, narrow passages create occlusions and limited visibility of the ground, causing gaps in both feature tracking and surface mapping. The presence of large protruding structures that appear ground-like in texture can lead to incorrect segmentation of the actual ground surface. These issues collectively challenge our EgoExo++ pipeline, occasionally resulting in an unreliable representation of the terrain.
VI Conclusion and Future Work
This work presents an AR-based framework to synthesize exocentric camera views from egocentric feed in real-time for improved underwater ROV teleoperation. A pose geometry-based closed-form solution is formulated for the proposed EgoExo++ problem and then integrated into a visual SLAM backbone. The end-to-end pipeline only requires a sequence of past egocentric views to generate 2D/2.5D exocentric views with the accurate ROV model projected onto them. The proof-of-concept is validated by ground plane estimation and reprojection error analyses in a series of 2D indoor navigation experiments. Subsequent field experiments are conducted to demonstrate the effectiveness of 2.5D scene rendering in unstructured underwater cave scenarios. A subjective study proves the advantage of the proposed TeleOp console over traditional systems and demonstrates that the framework: (i) offers more informative peripheral views, (ii) provides better situational awareness, and (iii) facilitates an interactive ROV teleoperation experience. We are currently exploring more comprehensive multi-sensor fusion-based underwater SLAM backbones, such as the SVIn2 [69], for more accurate and robust estimation. We further plan to develop and integrate more interactive features that would serve as a simulation platform for ROV teleoperation research.
Acknowledgements
This work is supported in part by the NSF grants , , and . The authors would like to acknowledge the help from Woodville Karst Plain Project (WKPP), El Centro Investigador del Sistema Acuífero de Quintana Roo A.C. (CINDAQ), Global Underwater Explorers (GUE), Ricardo Constantino, and Project Baseline in providing access to challenging underwater caves.